# StimuVAR: Spatiotemporal Stimuli-aware Video Affective Reasoning with Multimodal Large Language Models

### v1

@@ 4.3 Ablation Study
In this section, we look into how the proposed mechanisms affect StimuVAR. We conduct experiments on the VCE dataset [13].
@@
Valley V1 [21] 256 44.5 68.4
Valley V2 [21] 256 46.6 71.2
Valley V3 [21] 256 47.5 72.9
TS Selection [52] 128 47.8 72.5
@@
Figure 6: Comparison of tube score maps and output responses of different temporal aggregation designs. Valley [21] often hallucinates objects that do not appear in a video. TS Selection [52] could be easily misled by its inaccurate temporal selection. In contrast, our emotion-triggered tube selection simultaneously considers both spatial and temporal information, resulting in the accurate selection of emotion-relevant tokens. Relevant and irrelevant words are colored.
@@
Temporal aggregation design. Due to the computational cost of attention increasing with increasing numbers of visual tokens, it is challenging for MLLMs to encode videos effectively and efficiently. Valley [21] introduces three variants of temporal aggregation to encode video representations: (V1) averages the <Patch> tokens; (V2) computes the weighted sum of the <Patch> tokens; and (V3) combines the averaged <Patch> tokens with the dynamic information extracted by an additional transformer layer on the <Patch> tokens. Wang et al. [52] design a token selection module for efficient temporal aggregation, which considers temporal and spatial dimensions sequentially to select tokens; we name it TS Selection. The proposed StimuVAR uses emotion-triggered tube selection for temporal aggregation, considering spatiotemporal information simultaneously. In this experiment, we replace StimuVAR’s tube selection with these baseline temporal aggregation designs to compare their effectiveness. As shown in Table 4 (Right), StimuVAR’s emotion-triggered tube selection outperforms Valley V1 - V3. This is because Valley V1 - V3 fuse spatial features in the temporal dimension, which is prone to producing objects not appearing in the video in the feature domain, causing visual hallucinations. StimuVAR also outperforms TS selection, demonstrating the superiority of the spatiotemporal design of our tube selection. Fig. 6 illustrates the hallucination issue of Valley V1 - V3. Specifically, the fused features incorrectly produced objects like ’fozen body of water’, ’frozen lake’ and ’waterfall’, which are not present in the video. On the other hand, TS Selection relies heavily on temporal selection. As shown in the middle of Fig. 6, if frames containing emotion-related events are mistakenly filtered out by temporal selection, the final output would be misled. In contrast, StimuVAR uses tubes as the minimal unit of selection, mitigating such issues. Our method achieves higher performance while reducing the number of tokens to enhance efficiency.
Visualization of emotion-triggered tube selection. As discussed, the emotion-triggered tube selection strategy enables StimuVAR to focus on spatiotemporal stimuli regions while reducing computational costs by using fewer tokens. In Fig. 7, we present two more examples to visualize the token score maps. In the first example, higher scores are allocated to the spatiotemporal regions containing “a man standing on a rocky shore and a large water splash or explosion”, while lower scores are given to regions with less emotionally relevant information, such as the last frame. In the second example, the primary focus is on an unexpected event in the video where a turkey is chasing a man in a black outfit. Most remaining areas do not explicitly trigger viewers’ emotions and thus receive lower scores. Both examples demonstrate StimuVAR’s ability to select the essential tokens most likely representing emotional stimuli in the spatiotemporal domain. The visualization also shows a close alignment between the selected tubes and the keywords in the output explanations, highlighting how the proposed strategy benefits affective reasoning. Moreover, we observe that the emotional stimuli usually occupy only small portions of a video, indicating that many tokens are redundant to the VAR task. This explains StimuVAR’s high token efficiency, as shown in Fig. 8.
Figure 7: Visualization of StimuVAR’s tube score maps and corresponding output responses. From top to bottom rows: original frames, tube score maps, and tube score maps overlaid on the frames. Emotion-triggered tube selection identifies the most emotion-relevant spatiotemporal areas, enabling effective and efficient video representations for VAR. Relevant and irrelevant words are colored.
@@
Number of selections in stimuli awareness. We explore varying numbers of frames to sample and tubes to select in our event-driven frame sampling and emotion-triggered tube selection strategies, respectively. As shown in Fig. 8, initially increasing the numbers of selected frames $N$ and tubes $K$ improves performance. Specifically, the Top-3 accuracy increases from 69.9% to 73.5% with three more frames (from $N=3$ to $N=6$), and there is an improvement from 71.8% to 73.5% when we select the top $K=4$ tubes compared to $K=1$. These increases are attributed to the additional information available, which allows the model to gain a more comprehensive affective understanding of the videos. However, when we continue to increase the number of frames and tubes, the performance begins to decline. The Top-3 accuracy decreases to 72.9% when $N=12$ frames are sampled, and it also drops to 71.0% when we double the tubes to $K=8$. This decline is due to redundant information from excessive frames or tokens, which distracts the model from emotional stimuli. Still, StimuVAR consistently outperforms the MLLM baselines (as presented in Table 1) with these different $N$ and $K$, demonstrating that the proposed strategies are not very sensitive to hyperparameter choices or computational budgets.
@@
Figure 8: Ablation on the number of selected frames (Left) and tubes (Right). Results are on the VCE dataset.
@@
Property of event-driven frame sampling. As discussed in Section 3.1.1, the motivation behind proposing the event-driven frame sampling strategy is to effectively capture rapid, unexpected events that are most likely to evoke viewers’ emotional reactions. Table 4 (Left) validates the effectiveness of this strategy. In this experiment, we further analyze the properties of this strategy. Specifically, we notice that surprise is one of the emotions particularly sensitive to rapid, unexpected events. In other words, such events inherently tend to evoke an emotion of surprise. To explore this, we examine the models’ performance on the surprise test subset of the VCE dataset, which consists of 2,529 videos annotated as surprise. In LABEL:tab:Suprise, given the same number of frames to select (i.e., $N=6$), StimuVAR with event-driven frame sampling achieves 72.9% Top-3 accuracy on the surprise subset, while the variant that uses uniform sampling drops to 67.6%. Additionally, we evaluate Chat-UniVi [24], the MLLM baseline with the best performance in Table 1. With its original uniform sampling design, Chat-UniVi achieves only 18.1% on the surprise subset, significantly lower than its 38.6% overall accuracy on the entire test set. However, when equipped with event-driven frame sampling, its accuracy improves to 25.7% (+7.6%) on the surprise subset and to 41.0% (+2.4%) on the entire test set, which is gained by only changing the input frames. These results again demonstrate the effectiveness of event-driven frame sampling in capturing rapid, unexpected events as emotional stimuli at the frame level, leading to higher performance, especially for emotions sensitive to such events.
Figure 9: A failure case due to subjectivity. Although StimuVAR’s emotional prediction (“admiration”) differs from the ground-truth label (“satisfaction”), it gives a plausible explanation. This “incorrect” prediction arises from the inherent subjectivity of emotion tasks, emphasizing the importance of incorporating reasoning.
Failure cases due to subjectivity. Emotion prediction is a task with an extent of subjectivity. Although humans share certain commonsense knowledge, the same visual stimuli may still evoke different emotional responses in different viewers. We observe this interesting phenomenon from our results as well, where StimuVAR makes predictions that differ from the ground-truth labels, yet its rationales are reasonable. For example, in Fig. 9, StimuVAR correctly describes a woman working out in a gym. The woman’s dedication and commitment could evoke emotions of the predicted admiration, the ground-truth satisfaction, or both. These emotions are all reasonable. This example highlights the crucial role of reasoning, especially given the subjective nature of emotion tasks, as reasoning explains how the model makes its predictions and, consequently, how these results can be applied.
Table 6: Ablation on the generalizability across different LLMs on the VCE dataset.
       LLM        Llama2-7b [42]        Llama2-13b [42]        Llama3-8b [43]
       Top-3        73.5        73.6        74.0
Generalizability across different LLMs. We apply StimuVAR to the MLLM backbone with different LLMs. Table 6 shows that increasing the model size leads to performance improvement. Using the more advanced LLM, Llama3-8b [43], further improves performance to 74.0%. These results demonstrate that the proposed strategies have high generalizability across different backbone architectures and can consistently scale with larger/better backbones.
Extension to image data. Although StimuVAR is designed to predict viewers’ emotional reactions to videos, it can be extended to handle image inputs as well. This can be done by treating an image as a single-frame video. In this case, frame sampling is not useful, but token selection and affective training are still valid. We use EmoSet [18], a large-scale emotion image dataset, to evaluate StimuVAR’s performance on image data. EmoSet consists of 118,102 images manually labeled into one of eight emotion categories. In addition to prediction accuracy, we also measure the reasoning quality. To test StimuVAR’s direct generalizability from videos to images, we evaluate the model trained with VCE data without fine-tuning on Emoset. We compare StimuVAR with image-oriented MLLMs [44, 39, 40] and EmoVIT [17], where EmoVIT is explicitly fine-tuned on EmoSet. Table 7 shows that StimuVAR outperforms image-oriented MLLMs in both prediction accuracy and reason quality by decent margins. Notably, while EmoVIT, fine-tuned on EmoSet, achieves the highest prediction accuracy, its reasoning ability is limited. In the LLM-as-a-judge evaluation, StimuVAR’s reasoning wins in most samples, as EmoVIT fails to produce plausible reasoning consistently. These results demonstrate StimuVAR’s deep understanding of human emotions and its generalizability from videos to images.
Table 7: Quantitative comparison on the EmoSet dataset. EmoVIT* indicates that EmoVIT is fine-tuned on the EmoSet data, while the other methods are not.
Method Acc Emo-ali RR RW WR WW CLIP-S Win Lose Tie Score
BLIP-2 [44] 46.8 43.2 36.5 10.3 6.7 46.5 65.5 98 0 2 1.26
LLaVA [39] 44.0 40.1 35.3 8.7 4.8 51.2 66.4 42 13 45 3.06
InstructBLIP [40] 42.2 43.9 29.9 12.3 14.0 43.8 60.9 34 28 38 3.17
EmoVIT* [17] 83.4 26.7 24.2 59.2 2.5 14.1 62.8 68 9 23 2.33
StimuVAR (Ours) 67.7 68.5 66.3 1.4 2.2 30.1 70.5 - - - 3.34
5 Conclusion
In this work, we propose StimuVAR, an MLLM-based spatiotemporal stimuli-aware framework for VAR. The event-driven frame sampling strategy employs optical flow as a cue to sample frames of key events at the frame level, while the emotion-triggered tube selection strategy further localizes the emotional stimuli areas at the token level. Additionally, we create VAR visual instruction data used for affective training, enhancing StimuVAR’s affective understanding and reasoning ability. To the best of our knowledge, StimuVAR is the first MLLM-based method for VAR from the viewers’ perspective. Extensive experiments demonstrate that StimuVAR effectively and efficiently predicts viewers’ emotional reactions while also reasoning about its predictions. This dual capability enhances the model’s reliability, making it suitable for real-world applications.
Data availability statements. All training/testing data and code that support the findings of this study have been deposited in VCEVE-8 and EmoSet.

### v2

@@ 4.3 Ablation Study
In this section, we look into how the proposed mechanisms affect StimuVAR. We conduct experiments on the VCE dataset [mazeika2022would].
@@
Valley V1 [luo2023valley] 256 44.5 68.4
Valley V2 [luo2023valley] 256 46.6 71.2
Valley V3 [luo2023valley] 256 47.5 72.9
TS Selection [wang2022efficient] 128 47.8 72.5
@@
Figure 6: Comparison of tube score maps and output responses of different temporal aggregation designs. Valley [luo2023valley] often hallucinates objects that do not appear in a video. TS Selection [wang2022efficient] could be easily misled by its inaccurate temporal selection. In contrast, our emotion-triggered tube selection simultaneously considers both spatial and temporal information, resulting in the accurate selection of emotion-relevant tokens. Relevant and irrelevant words are colored.
@@
Temporal aggregation design. Due to the computational cost of attention increasing with increasing numbers of visual tokens, it is challenging for MLLMs to encode videos effectively and efficiently. Valley [luo2023valley] introduces three variants of temporal aggregation to encode video representations: (V1) averages the <Patch> tokens; (V2) computes the weighted sum of the <Patch> tokens; and (V3) combines the averaged <Patch> tokens with the dynamic information extracted by an additional transformer layer on the <Patch> tokens. Wang et al. [wang2022efficient] design a token selection module for efficient temporal aggregation, which considers temporal and spatial dimensions sequentially to select tokens; we name it TS Selection. The proposed StimuVAR uses emotion-triggered tube selection for temporal aggregation, considering spatiotemporal information simultaneously. In this experiment, we replace StimuVAR’s tube selection with these baseline temporal aggregation designs to compare their effectiveness. As shown in Table 5 (Right), StimuVAR’s emotion-triggered tube selection outperforms Valley V1 - V3. This is because Valley V1 - V3 fuse spatial features in the temporal dimension, which is prone to producing objects not appearing in the video in the feature domain, causing visual hallucinations. StimuVAR also outperforms TS selection, demonstrating the superiority of the spatiotemporal design of our tube selection. Fig. 6 illustrates the hallucination issue of Valley V1 - V3. Specifically, the fused features incorrectly produced objects like ‘fozen body of water’, ‘frozen lake’ and ‘waterfall’, which are not present in the video. On the other hand, TS Selection relies heavily on temporal selection. As shown in the middle of Fig. 6, if frames containing emotion-related events are mistakenly filtered out by temporal selection, the final output would be misled. In contrast, StimuVAR uses tubes as the minimal unit of selection, mitigating such issues. Our method achieves higher performance while reducing the number of tokens to enhance efficiency.
Fine-tune state-of-the-art video-oriented MLLMs using VAR instruction data. To exclusively assess the contributions of our proposed VAR instruction data, we use our VAR instruction data to fine-tune the three best-performing MLLMs in Table 1: Chat-UniVi [Jin_2024_CVPR], Valley [luo2023valley], and VideoChat2 [li2024mvbench], where we follow their official fine-tuning guidelines for each model. We evaluate them on the VCE, VE-8 and YF-6 datasets, using the same evaluation metrics described in Sec. 4.1. As can be seen in Table 6, fine-tuning using our VAR instruction data consistently improves the performance of all three MLLMs on all three datasets. This highlights the quality and effectiveness of the proposed VAR instruction data in enhancing MLLMs’ affective reasoning capabilities. On the other hand, compared to these fine-tuned MLLMs, our StimuVAR still consistently demonstrates clear superiority. This highlights the advantages of the proposed model design. Overall, these results demonstrate that both the proposed VAR instruction data and model design are effective, playing their roles in contributing to excellent performance.
Table 6: Quantitative comparison of state-of-the-art video-oriented MLLMs, with and without fine-tuned on our proposed VAR instruction data.
Methods Acc Emo-ali RR RW WR WW CLIP-S Win Lose Tie Score
VCE
Valley [Uncaptioned image] [luo2023valley] 31.3 29.4 19.2 12.1 10.2 58.5 69.4 45 12 43 2.89
Valley [Uncaptioned image] [luo2023valley] 42.1 43.0 38.4 3.7 4.6 53.3 70.5 32 14 54 3.21
VideoChat2[Uncaptioned image] [li2024mvbench] 31.1 36.4 24.0 7.1 12.4 56.5 68.6 56 12 32 2.65
VideoChat2[Uncaptioned image] [li2024mvbench] 40.1 40.2 28.6 11.5 11.6 48.3 69.4 39 16 45 3.15
Chat-UniVi[Uncaptioned image] [Jin_2024_CVPR] 38.6 29.5 21.0 17.6 8.5 52.9 70.2 50 19 31 2.83
Chat-UniVi[Uncaptioned image] [Jin_2024_CVPR] 56.8 48.5 45.3 11.5 3.2 40.0 71.8 36 15 48 3.22
StimuVAR (Ours) 73.5 69.6 68.8 4.7 0.8 25.6 75.3 - - - 3.40
VE-8
Valley [Uncaptioned image] [luo2023valley] 16.1 20.6 12.2 4.0 8.5 75.3 67.0 35 7 58 2.88
Valley [Uncaptioned image] [luo2023valley] 25.5 26.6 21.1 4.4 5.5 69.0 68.4 33 17 50 3.05
VideoChat2[Uncaptioned image] [li2024mvbench] 16.5 15.6 14.3 2.2 1.4 82.1 65.1 48 13 39 3.02
VideoChat2[Uncaptioned image] [li2024mvbench] 32.7 29.9 27.1 5.6 2.8 64.5 67.9 27 22 51 3.12
Chat-UniVi[Uncaptioned image] [Jin_2024_CVPR] 14.4 19.2 9.9 4.5 9.3 76.3 65.9 37 20 43 3.08
Chat-UniVi[Uncaptioned image] [Jin_2024_CVPR] 25.4 28.6 20.1 5.3 8.5 66.1 66.3 26 19 55 3.15
StimuVAR (Ours) 40.5 41.7 33.7 6.8 8.0 51.5 70.6 - - - 3.29
YF-6
Valley [Uncaptioned image] [luo2023valley] 10.0 18.9 5.9 4.1 13.0 77.0 66.4 44 10 46 2.89
Valley [Uncaptioned image] [luo2023valley] 18.0 22.0 12.6 5.4 9.4 72.6 66.8 28 19 53 3.14
VideoChat2[Uncaptioned image] [li2024mvbench] 24.3 23.7 18.5 5.8 5.2 70.5 64.1 37 16 47 3.03
VideoChat2[Uncaptioned image] [li2024mvbench] 39.7 35.6 34.1 5.6 1.5 58.8 65.1 30 21 49 3.16
Chat-UniVi[Uncaptioned image] [Jin_2024_CVPR] 28.4 29.2 22.1 6.3 7.1 64.5 66.3 33 14 53 3.12
Chat-UniVi[Uncaptioned image] [Jin_2024_CVPR] 32.7 30.3 22.8 9.9 7.5 59.8 67.0 26 22 52 3.18
StimuVAR (Ours) 46.2 44.2 37.3 8.9 6.9 46.9 72.1 - - - 3.30
@@
Number of selections in stimuli awareness. We explore varying numbers of frames to sample and tubes to select in our event-driven frame sampling and emotion-triggered tube selection strategies, respectively. As shown in Fig. 7, initially increasing the numbers of selected frames $N$ and tubes $K$ improves performance. Specifically, the Top-3 accuracy increases from 69.9% to 73.5% with three more frames (from $N=3$ to $N=6$), and there is an improvement from 71.8% to 73.5% when we select the top $K=4$ tubes compared to $K=1$. These increases are attributed to the additional information available, which allows the model to gain a more comprehensive affective understanding of the videos. However, when we continue to increase the number of frames and tubes, the performance begins to decline. The Top-3 accuracy decreases to 72.9% when $N=12$ frames are sampled, and it also drops to 71.0% when we double the tubes to $K=8$. This decline is due to redundant information from excessive frames or tokens, which distracts the model from emotional stimuli. Still, StimuVAR consistently outperforms the MLLM baselines (as presented in Table 1) with these different $N$ and $K$, demonstrating that the proposed strategies are not very sensitive to hyperparameter choices or computational budgets. Regarding computational cost, as expected, increasing the number of selected frames $N$ and tubes $K$ results in longer inference times (about linearly). This highlights the importance and effectiveness of the proposed event-driven frame sampling and emotion-triggered tube selection strategies, which can improve both accuracy and computational efficiency.
@@
Figure 7: Ablation on the number of selected frames (left) and tubes (right) for accuracy and inference time on the VCE dataset.